OpenStudy (anonymous):

If A and B are events, P(A)= .3, P(A or B)= .7, and P(A and B)= .05, what is P(B)?

OpenStudy (nikvist):

\[P(A\cup B)=P(A)+P(B)-P(A\cap B)\] \[P(B)=P(A\cup B)+P(A\cap B)-P(A)=0.7+0.05-0.3=0.45\]